OpenStudy (anonymous):

How do you: Use f(x) = 3x -5 and g(x) = 2 - x^2 to evaluate the following expressions: 1. (a) f(g(0)) (b) g(f(0)) 2. (a) (f ` g) (-2) (b) g(g(3))

OpenStudy (anonymous):

I will do one of them as an example for you, tell me which one?

OpenStudy (anonymous):

1 b

OpenStudy (anonymous):

For 1.b part, you need to find f(0) first. f(x) = 3x - 5 f(0) = -5 \[g(x) = 2-x^2 = g(fx) = 2 - (f(x))^2\] g(f(0)) = g(-5) = \(2 - (-5)^2 = 2 - 25 = -23\)

OpenStudy (anonymous):

why does f(0)= -5?

OpenStudy (anonymous):

f(x) = 3x - 5 f(0) = 3*0 - 5